Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Enter a number. Round your answer to four decimal places.)

μ = 26; σ = 3.6

P(x ≥ 30) =
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

P(X>=30)

= P(Z> = (30-26)/3.6)

= P(Z> = (1.11))

= .1333